EVA Report November 9th

EVA Report Sol 6

The EVA started at 10:50 AM today. Knowing that Brazilian TV was coming, I didn’t know for sure when they’d be here and ready to film us outside. They also needed to interview a few of the crewmembers on their projects before we could start. They wanted some footage of us on the Rover and ATVs, which we drove about 100 meters past the MDRS sign and back to the Hab area. I understand that I did not request the use of the rovers for today’s EVA, but I didn’t know they’d want footage of us on the rovers until this morning. Something to note was that the red ATV had some trouble starting compared to yesterday. I had to pump the throttle in order to get it going.

After parking the rovers, we collected our samples from Sample Location #1, the first location we sampled from on EVA #1. Once we got the sample, we walked into the small gully just north of the Hab. We took samples from 3 different locations along the hillside west of the hab. The “popcorn” like surface of the regolith can be rather unstable, as I found out the hard way (by falling). Nothing broke, I wasn’t injured, the helmet was intact, and the fan was still working, so luckily I survived! After photographing the area, we returned to the hab and recompressed by 12:45 PM.

In Hindsight, I’ll have to be more careful on even very shallow slopes, as it doesn’t take much too loose traction, even when wearing boots. To make sure no one else falls by making the same mistake, I’ll have to brief the crewmembers on how to lower their center of gravity (by squatting) when ascending or descending hills. I did not have time to brief the crew members on radio Comms voice procedures since we had the Brazilian TV crew here today. This will have to be done before tomorrow’s EVA.

Commander Report November 9th

My soundtrack of last night: ACDC live in River Plate stadium, Buenos Aires.:) Camila did a delicious fried rice. Thank You!:)

4h30m – I woke up early today. Sometimes the heating system switches off. So we need to check. This is also a possibility in Mars. Many times, in Mars, the heating systems will not function properly. So, will be necessary a lot of resilience about. Try to fix or deal with the consequences.

The commander’s room is unique with natural lighting. This helps me to woke up early.

7h02m – I walked 40 minutes, something around 2,5 km. Was 1 degreeC. In the beginning of the exercise was very difficult. I never did exercises in a temperature like this. This is also about resilience. In Mars the temperature can be so low. According to Space.com, on average, the temperature on Mars is about minus 80 degrees F (minus 60 degrees C). In winter, near the poles temperatures can get down to minus 195 degrees F (minus 125 degrees C). A summer day on Mars may get up to 70 degrees F (20 degrees C) near the equator, but at night the temperature can plummet to about minus 100 degrees F (minus 73 C). (Source: https://www.space.com/16907-what-is-the-temperature-of-mars.html)

Something very pleasant today was watch the sunrise in the mountains. I perceive how We are not aware about how the cosmic influences are affecting life. Just look the sun! It is a star and it is so close and far from us, influencing so many cycles of our life, from the agriculture to our communication systems.

8h40m – The Brazilian TV Globo arrived earlier than We imagined. From the MDRS corridors I talked with Luigi Sofio (reporter) and Lucio Rodrigues (cinegraphist) and introduced myself. Was good to talk with someone on portuguese. Since their arrival until 13h, We presented all the attention to Them. They did images of our systems and procedures and also the research activities.

11h – We did the EVA activity. Was very hard. Was my first EVA. Would possible to see how the spacesuit would restricts simple human activities. I felt excited with it. This moves me. A spacesuit would be my future?

We did a lot of work today, outside. Today was very shiny and I considered that Brazilian TV did very, very good images. What sincronicity We have a bright day as today.

In the EVA We collected soil samples and visit the ravines and some mountains of Mars. Was hard climb some mountains wearing the spacesuits, considering the limitations of movement. It is necessary presents a great attention and care to avoid an accident. But the EVA was a challenger and an achievement.

13h – I prepared the lunch: small bakes of rice and cheddar with broccoli and peas.

13h40m – We lunched. We had a lot to clean up.

15h40m – The guys of the EVA: Me (Julio), Brandon and Atila felt tired, because the sun radiation was to high today. So everybody decide take a nap.

17h40m – I woke up from my nap. It is good remember about the CapComm later. Good time to listen Slayer and finish the reports. J

18h20m – Camila and Carmen prepared sweet rice for tonight.

19h – CapComm. Today I walked the total of 5,8 km.

Sol Summary November 9th

MDRS Daily Summary Report for sol 0.6

Summary Title: A shine and hot day.

Mission Status: The internet is so low today. The day circles around Globo TV visit. They arrived in MDRS at 8h40m and leaved 13h. We support them in interviews, images and guidelines.

Sol Activity Summary: In the EVA today We walked more, compared to other day and also was so shiny.

Look Ahead Plan: Tomorrow, We are planning an EVA in the morning to study some plants around MDRS to the research of Camila. Julio Rezende will work in his book. Atila and Carmen will develop their research also.

Anomalies in work: Nope.

Weather: 9o C, 49o F

Crew Physical Status: Fine.

EVA: today happened EVA with the ATV hide and walking in the ravines and Mars mountains.

Reports to be file: We sent in CapComm.

Support Requested: None.

Science Report November 9th

Mars has been also great today! We had some sunshine, but it wasn’t as warm as other days. In fact, I had problems finding some sunshine for my seeds. But alright, continuing with the Science report:

The autoclave is already set and working! We read carefully the manual to use this, because I’m particularlly fearful of that machine. But I noticed that it requires a great amount of destillated water for running. I would like to ask about this issue because I had to use a gallon and a half for one sterelization cycle. I’ll need to do another one in the next week, so I would like to know if it is possible to get more gallons of destillated water.

We also need to figure out the position of the machines at the lab. The autoclave requires a good table, so it occupies the only lab table at the lab. I will not be using the hood for my project, so I need to move it somewhere else. Right now, the lab is pretty messy.

As for the projects, there has been some soil sampling today during the EVAs. These samples have been collected for both typing and using during the germination experiments.

No photos are going to be attached to this report, because my main concern is with the destillated water.

Greenhab Report November 6th

On Sol 3 the Green Hab (GH) was quite nice: I checked the humidity 3 times on the day, every time it has 16-17% (tomorrow I’m going to be more specific with the times). I opened the door of the GH at 10:45 and I closed it at 17:30, it was open because it’s getting pretty hot inside. The temperature inside the GH fluctuated between 25°C-32°C during the day.

Inside the GH there is some seed from a former experiment someone was doing, Shannon told me to water them. I did 2 times today at 10:45 and 15:40, I was about to do a third at 18:20 but they had enough water, I will try to water them tomorrow before 09:00.

Species (common name) Number of bowls
Tomatoes 25
Cucumber 7
Melon 8
Beans 17
Pepper 9

At Sol 4 I’ll have to re-label most of the bowls because it’s a mess. So I’ll let you know how that worked. Also, most of the GH is a mess, so as we doing in the lab I’ll try to rearranged everything. It’s a pretty nice place and I think it can be very useful. We found a red wheelbarrow that we are trying to build up to use it at the EVA’s, but it’s extremely weird, we don’t know if it has missing parts or what, but we can’t finish to build it up. It’s driving me crazy.

Tomorrow I’ll start my project, I’m going to measure the rate of germination of quinoa, amaranthus, and mustard.

Sol Summary November 4th

We came in safe and sound, landing went smooth without incident.We had the opportunity to observe the remains of the famous Martian Dinosaurs that dot the landscape. We will have to wait until tomorrow to send out the first report, as we are attending the traditional martian bonfire to start off the field season.

Greenhab Report November 7th

Atila Meszaros

Sol 4 – 7/11/2017

Ambient (no heating or/cooling): None

Ambient with window/door open: Door opened at 09:10 and closed at 17:15

Heating: None

Cooling: None

Shade cloth on/off: Off

Changes to crops: None

Daily water usage for crops: 2 gal approximately

Time(s) of watering for crops: 08:54 and 13:30

Changes to research plants: Atila’s project just started today. Tomorrow I will send more feedback.

Daily watering and amount of water used: none

Narrative:

Hi Key! Atila in here. Sol 4 was a very busy day at the Green Hab (GH).

Ok, so let’s start with some crazy data about the GH today:

Time 08:54 11:52 13:30 17:15
Temperature inside (°C) 29 35.5 38 14
Temperature outside (°C) 5 13 14 8
% Humidity 16 16 16 21

*The times in red were when I watered the seed that is still in the lab, they are supposed to be someone else experiment.

Shannon told me to water the seeds 3 times every day, but by 18:00 they were pretty wet, I don’t want them to rot.

On the noon it was extremely hot, jeez! I had to take my shirt off for a while. I was alone, so I guess it wasn’t that weird.

As I told Sylvain on Sol 3, today I re-labeled all the seeds from the former experiment, it was a mess. Now the seeds look great, I didn’t group them, so they still have the random distribution.

I didn’t have enough time to rearrange the GH because I started my experiment today. I’m going to measure the germination percentage and rate of quinoa, amaranthus and two kinds of mustard on a control soil and the mars analog soil. I’ll attach you a picture of the pots. I still have to finish planting the seed I’m going to put on the growth chamber that’s in the Lab. I’m aiming to finish it before dinner.

Also, I’m aiming to make the GH great again by tomorrow afternoon, I’ll let you know how that worked.


Operations Report November 7th

SOL: 4

Crew Carmen Atauconcha

Generator— It is working perfectly. On.

Solar— Working perfectly

Diesel – 70%

Propane – 57psi

Water (trailer) –

Water (static) – 15cc from top

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter:

ATV’s Used: No

Oil Checked: No

Atv # Fuel Used Gals: Not used

Tires Status:

# Hours Used Day: No EVA’s today

Phobos used: No

Hours:0

Beginning charge:-

Ending charge:-

Currently charging: –

Deimos used: No

Hours:-

Beginning charge:-

Ending charge: –

Currently charging:-

HabCar used? No

Notes and Comments:

Today we had a meeting, where we decide how to reduce the usage of water. Also, we review the radios again and separate which ones are working. In some of them the sound is low. For tomorrow we are planning to organize and clean the tools and equipments that we have in the Hab.

Summary of operations activities:

Everyone is working in their researches. For tomorrow we are going to have an EVA session.

Questions and Concerns to Mission Support:

Hello mission support, I have a question about phobos and deimos, Which are the deimos, and which are the phobos? Please.

Sol Summary November 7th

MDRS Daily Summary Report for sol 4

Summary Title: A preparation day.

Mission Status: The best thing to do is to try to keep a good relationship and a High State of Consciousness in our crew. We are going fine in our IVA and EVA. We must keep this.

Sol Activity Summary: The commander reflected about sustainability dimensions to be observed in MDRS and also in possible space travel and colonies in Mars.

Look Ahead Plan: Our plans today was to be in the MDRS and start and support the agriculture research developed by Atila. Tomorrow, We you do an EVA. Our EVA officer Will present more details. The commander still working in his book related to Mars and Sustainability research.

Anomalies in work: Nope. Just becoming very cold. 😞

Weather: 8 C.

Crew Physical Status: Fine.

EVA: No EVA today.

Reports to be file: We sent in CapComm.

Support Requested: None.